OpenStudy (anonymous):

Given the function g(x) = 8x − 2, find g(−2).

OpenStudy (lgbasallote):

substitute -2 to every x of g(x) 8(-2) - 2 = -16 - 2 = -18

OpenStudy (anonymous):

THANKYOU!!!!!!